A child receives a balloon filled with 2.42 L of helium from a vendor at an amusement park. The temperature outside is 311 K. When the child brings the balloon into an air-conditioned room, the volume of the balloon decreases to 2.30 L. What is the temperature of the air-conditioned house? Assume the pressure in the balloon stays the same.


1
Expert's answer
2021-05-24T07:19:08-0400

Here, "V_1=2.42L,T_1=311K,V_2=2.3L,T_2=?"


As The pressure is constant,

Using charle's law we have-


"V_1T_2=V_2T_1\\[9pt]\Rightarrow2.42\times 311=2.30\times T_2\\\\[9pt]\Rightarrow T_2=\dfrac{2.42\times 311}{2.30}=327.22K"